# What happens when generative AI models train recursively on each others’ generated outputs?

### v2

@@ 6 Experiments with Large Language Models
Setup. To produce the initial set of models, we fine-tune $K=3$ instances of Facebook’s OPT-350m [70] on task-specific datasets to simulate the effect of training on $(D_{*},\tilde{D}_{k})$ at time $t=0$. We use OPT-350m as the base model because its training data is publicly available, allowing us to know the contents of $D_{*}=$ BookCorpus [67], CC-Stories [12], the English portion of CommonCrawl, and public Reddit data. For our experiment, we approximate $D^{*}$ using only BookCorpus due to practical constraints. Our task specific datasets $\tilde{D}_{k}$ are SciQ [39], a dataset of physics, chemistry, and biology questions, OpenAI’s GSM8K, a dataset of grade school level math problems [20], and ai2_arc [18], a dataset of grade school science questions, chosen to given each model a distinct private task. The goal of each dataset is to train models that answer questions in the dataset domain with high accuracy and fluency, and we fine-tune one model on each dataset.
After each generation, we use $\hat{\theta}_{tk}$ to produce synthetic data $D_{t+1,k}$ for use in the next update. To construct $D_{t+1}$, we randomly sample prompts from $\tilde{D}_{k}$ for each $k$ and feed them to $\hat{\theta}_{tk}$, which completes the text. We use $\alpha\in\{0,0.5,1\}$ and $\beta\in\{0,0.5,1\}$ for $T=15$ generations of training. For the $t$-th training generation, model parameter $\hat{\theta}_{tk}$ is obtained by finetuning on a dataset of size $n=12,500$ drawn i.i.d. from the datasets $\tilde{D}_{k}$, $D_{*}$, $D_{t}$ with weights $\bar{\alpha}\beta$, $\bar{\alpha}\bar{\beta}$, and $\alpha/K$, respectively.
@@
Figure 3: Predicted behavior over time for a $K=3$ model system with varying $\alpha$, $\beta$. We use equations for MSE from theorem 2 and run simulations with $K=3$, dimension 50, rank 15.
Figure 4: Actual behavior over time for interactions between OPT models ($K=3$) with varying $\alpha,\beta$
Results. We show predicted and actual results of this experiment in Figures 5 and 6, respectively. Each figure plots the prediction error/loss of each model on its original task (first and third columns) and the other model’s task (second and fourth columns) to visualize how data-mediated interactions affect performance on each task. The theoretical results (Figure 5) show that a variety of behaviors are possible. Models sometimes get slightly worse or become better at certain tasks. As $\alpha\rightarrow 1$, the predicted error rises. These results are mirrored in the interactions of three OPT language models (Figure 4), and extend to interactions between two OPT models shown in Figure 6 in the appendix.
Of particular note is the experimentally observed homogeneity in model behavior on each task for $\alpha=0.5$ across all $\beta$s. In the $\alpha=0$ setting, each model performs worse at the other model’s task—e.g. model 1 has loss around 3 for task 2, while model 2 has loss around 2. However, once $\alpha>0$, the loss of model 1 drops to around 2 for task 2, matching that of the other model. This homogenization of task performance occurs for the other task as well, highlighting the benefits (improved task performance) and downsides (potential performance homogenization) of training on other models’ outputs. Additionally, when $\beta>0$, we see that learning information about task 1 helps task 3 because they are both science based questions, and vice versa.
@@ 7 Discussion
Limitations. Our work has several limitations. First, we outline objections to our claims about the increasing presence of generated outputs in training datasets (e.g. [25]) and counterarguments in Appendix B. Second, our theoretical framework assumes that new data in model updates is purely synthetic. In reality, if internet scrapes are used to create model update datasets, they will contain both synthetic and real data. Finally, we assume that model trainers use new scraped data for each model update but only reuse data from initial training. This assumption may limit the range of outcomes.
Broader Impacts. If data-mediated interactions homogenize generative models, causing them to coalesce on certain viewpoints, this could lead to pervasive bias in AI-generated content. Peterson [49] discusses this possibility, but focuses on the narrowing effect of AI rather than homogenization across models. Recent work from Wenger and Kenett showed homogeneity across creative outputs from many LLMs, suggesting these homogenization effects may already be felt by models [65]. Much future study is needed to evaluate the extent to which data-mediated interactions fuel homogeneity (as opposed to other causes) and develop mitigations.
Conclusions and Future Work. We provide a first look at possible outcomes of genAI models trained on each others’ data and find mixed effects. Training on other models’ data exposes models to concepts possibly missed in their own training data, but can homogenize model behaviors. Future work could consider additional nuances of interactions between models, explore how these interactions evolve in other modalities like image generation, and investigate whether fixed points (e.g. like the universal $\pi^{2}/6$ pathway of [23]) exist under this paradigm.
@@ Appendix A Exact training data statements from LLM papers
Publicly released on HuggingFace [13].
GPT4+ [7] No information provided No information provided
Grok 3 [66] No information provided No information provided
Jamba [59]
@@
[[[[Supervised fine tuning]]]] leverages highly
@@ Appendix B Counterarguments
Can’t we use watermarks to filter AI-generated content from future internet-scraped datasets? Several companies have publicly stated that they watermark AI-generated content [22, 10, 19], making this argument plausible. Furthermore, [25] show that using watermark detection techniques can help avoid model collapse under certain circumstances. However, reliance on watermarking has two major issues. First, watermarks are difficult to reliably detect and/or easily removed from generated content [68, 44, 54]. Second, detection via watermark requires sharing of watermark detection information, which is essentially a game-theoretic problem that relies on other companies’ willingness to cooperate. Both issues make watermarks an unreliable mechanism to rid datasets of AI-generated content.
@@
What if there’s less AI-generated online content than we think? Some work suggests there may be less AI-generated content online than previously postulated [47]. However, other works consistently point to an uptick in AI-generated content online [57]. We believe the widespread adoption and use of generative AI models across industries [8, 32, 53], particularly for use in content creation [1], provides strong evidence that AI-generated content will become a regular part of online life. Further empirical work is needed to vet both claims.
@@
What if one model provider dominates online content? Although numerous generative AI models are available online, some evidence suggests that one or two companies may dominate the AI landscape. A market research firm estimated that in January 2025, Open AI’s ChatGPT had 340 million monthly active users, Microsoft Copilot had 11 million, Google Gemini had 80 million, and Anthropic’s Claude had 2 million [72]. If only one model/company dominates, our paradigm of training on other models’ data will no longer be relevant and collapses back to the single model setting of prior work, e.g. [56, 41, 23]. Currently, though, this market research suggests that there are several models used by millions of users, making our assumptions somewhat reasonable. Future research could analyze market trends in model use and adoption to determine realistic assumptions.
@@
Appendix C Code for experiments
@@
Appendix D Proofs for results in Section 5
@@
D.1 Proof of Theorem 1
We now derive the distribution of $\hat{\bm{\theta}}_{t}$ conditional on the initial data $D_{0}$ given by $(\tilde{\bm{X}},\tilde{\bm{y}})$ and $(X_{*},y_{*})$. Specifically, we show that the estimates are Gaussian with mean and variance
@@ D.1 Proof of Theorem 1
For the inductive case, assume that the stated distribution holds up to generation $t-1$. From the definition of the workflow, (21) holds with $\bm{w}_{t}\sim\mathsf{N}(0,\sigma^{2}\mathrm{I})$ independent of everything else. Thus $\hat{\bm{\theta}}_{t}$ is Gaussian with mean
@@ D.2 Proof of Theorem 2
$\begin{bmatrix}\tilde{\bm{X}}^{+}\tilde{\bm{y}}\\ X_{*}^{+}y_{*}\end{bmatrix}\sim\mathsf{N}\mathopen{}\mathclose{{}\left(\begin{bmatrix}\tilde{\bm{S}}\tilde{\bm{S}}^{+}&0\\ 0&S_{*}S_{*}^{+}\end{bmatrix}(\bm{1}_{K+1}\otimes\theta),\sigma^{2}\begin{bmatrix}\tilde{\bm{S}}^{+}&0\\ 0&S^{+}_{*}\end{bmatrix}}\right).$ (26)
@@
Next, suppose that $\bm{G}_{1},\dots,\bm{G}_{t}$ are full rank and the stated distribution holds up to time $t-1$. By Theorem 1 we know that $\hat{\bm{\theta}}_{t}$ is Gaussian and so all that remains is to verify the given expressions for the mean and covariance. By the linearity of expectation and (21), the mean satisfies
@@ D.3 Proof of Theorem 3
D.4 Proof of Lemma 1
If $\alpha=0$ or if $\bm{S}=\bm{0}$ then $\bm{Q}=\bm{0}$ and so the stated result holds. Henceforth, we assume $10<\alpha<$ 1 and $\bm{S}$ is nonzero. Suppose that $\gamma\bm{S}=\lambda\tilde{\bm{S}}+(1-\lambda)(\mathrm{I}_{K}\otimes S_{*})$ for some $0<\beta\leq\lambda\leq 1$ and $\gamma>0$. Then,

### v3

@@ 6 Experimental Evaluation
Experiment setup. Training large language models from scratch is computationally infeasible for us, so we simulate the initial setup of language models trained on dataset $(D_{*},\tilde{D}_{k})$ at time $t=0$ by fine-tuning $K=2$ instances of a given pre-trained model architecture on carefully chosen $D_{*},\tilde{D}_{k}$. We experiment with two language model architectures—OPT-350m (Zhang et al., 2022) and Llama 3.2 1B (Dubey et al., 2024)—to assess how interactive training scales. Public sources (Touvron et al., 2023a; Zhang et al., 2022) state that both models were pretrained on BookCorpus (Yang, 2022), CC-Stories (Anderson, 2022), the English portion of CommonCrawl, and public Reddit data. We approximate $D_{*}$ with BookCorpus due to practical constraints. Each model is given its own initial task-specific dataset $\tilde{D}_{k}$ —SciQ (Johannes Welbl, 2017) (science questions) for the $k=0$ model; and OpenAI’s GSM8K (Cobbe et al., 2021) (grade school level math problems) for the $k=1$ model.
Interactive training proceeds as outlined in Figure 1, with fixed $\alpha$, $\beta$ values for each experiment. After training a new model generation $\hat{\theta}_{tk}$, we use $\hat{\theta}_{tk}$ to produce synthetic data $D_{t+1,k}$ that becomes part of the next generation’s training data (if $\alpha>0$). We produce $D_{t+1,k}$ by randomly sampling prompts from $\tilde{D}_{k}$ for each of the $K$ models and prompting $\hat{\theta}_{tk}$ to complete the text.
Training and evaluation. We run experiments on $K=2$ model systems with $\alpha\in\{0,0.5,1\}$ and $\beta\in\{0,0.5,1\}$, each for $T=15$ generations of training. At each training generation, models are fine-tuned on datasets of fixed size $n=12,500$ drawn i.i.d. from the datasets $\tilde{D}_{k}$, $D_{*}$, $D_{t}$ with weights $\bar{\alpha}\beta$, $\bar{\alpha}\bar{\beta}$, and $\alpha/K$, respectively. This effectively mimics the accumulate and subsample setup of Kazdan et al. (2025) with the additional wrinkle of data-mediated model interactions.
@@
Table 3: Change in loss behavior for $K=2$ interacting models at $\beta=0.5$. We show results as $\texttt{initial}\rightarrow\texttt{final}$ prediction loss values for models on their own and the other models’ tasks, at $\texttt{initial}=T=0$ and $\texttt{final}=T=15$ generations. For clarity, we colorize loss increase, decrease, and constancy ($\Delta\leq 0.1$).
(a) OPT models
$\alpha=0$ $\alpha=0.5$ $\alpha=1.0$
Model 1
on Task 1
$\rightarrow 3.3 3.33.3$$\rightarrow 3.3$ $3.3\rightarrow 3.5$
Model 2
on Task 2
$1.8\rightarrow 1.7$ $1.8\rightarrow 1.7$ $1.8\rightarrow 2.2$
Model 1
on Task 2
$3.1\rightarrow 3.5$ $3.1\rightarrow 1.8$ $3.1\rightarrow 2.2$
Model 2
on Task 1
$5.1\rightarrow 5.1$ $5.1\rightarrow 3.5$ $5.1\rightarrow 3.5$
(b) Llama 3.1 models
$\alpha=0$ $\alpha=0.5$ $\alpha=1.0$
Model 1
on Task 1
$2.8\rightarrow 2.9$ $2.8\rightarrow 2.9$ $2.8\rightarrow 3.0$
Model 2
on Task 2
$1.2\rightarrow 1.2$ $1.2\rightarrow 1.3$ $1.2\rightarrow 1.9$
Model 1
on Task 2
$2.0\rightarrow 2.5$ $2.0\rightarrow 1.4$ $2.0\rightarrow 1.8$
Model 2
on Task 1
$3.7\rightarrow 4.2$ $3.7\rightarrow 3.0$ $3.7\rightarrow 3.0$
Figure 3: PCA of embeddings of outputs produced by models $\theta_{t1}$ and $\theta_{t2}$ on datasets $\tilde{\mathcal{D}}_{1}$, $\tilde{\mathcal{D}}_{2}$, and $\tilde{\mathcal{D}}_{*}$. Results for $K=2$ Llama models trained with $\beta=0.5$ and varying $\alpha$. Leftmost plot shows embeddings at generation $t=0$, which are identical for all $\alpha$; while right plots show embeddings at $t=15$ with different $\alpha$.
Table 4: Cosine similarity of embedded output representations for $K=2$ Llama models at $\beta=0.5$. We show $\texttt{initial(t=0)}\rightarrow\texttt{final(t=15)}$ for $\tilde{\mathcal{D}}_{1}$ $\tilde{\mathcal{D}}_{2}$, and $\mathcal{D}_{*}$. We colorize increase and decrease.
$\alpha=0$ $\alpha=0.5$ $\alpha=1.0$
$\tilde{\mathcal{D}}_{1}$ $0.73\rightarrow 0.70$ $0.73\rightarrow 0.88$ $0.73\rightarrow 0.86$
$\tilde{\mathcal{D}}_{2}$ $8\rightarrow 0.79 0.880$$8\rightarrow 0.93$ $0.88\rightarrow 0.94$
$\mathcal{D}_{*}$} $0.50\rightarrow 0.59$ $0.50\rightarrow 0.64$ $0.50\rightarrow 0.58$
Results. As predicted in Figure 2, an $\alpha=0.5$, $\beta=0.5$ setting produces optimal results in terms of model performance across tasks. Table 3 reports the change in models’ loss values between the first and last ($t=15$) training generation for the $\beta=0.5$ setting. When $\alpha=0$, the same (human-generated) data is used for each training update, resulting in stable or slightly worse performance on different tasks, as models are forced into a local minima. When $\alpha=1.0$, models are only trained on generated outputs. They degrade on their original task due to the lack of real data but improve slightly on the other model’s task. However, at $\alpha=0.5$, models perform well on their own tasks and improve on the other model’s task. Results for other $\beta$ and $K=3$ are in Appendix and echo findings here.
While some amount of mixing improves model performance on previously-unseen tasks, homogenization occurs for $D_{*}$ at all $\alpha$ and for $\tilde{D}_{k}$ when $\alpha>0$—everywhere it can. Figure 3 shows the benefit of interactive training when $\alpha,\beta$ are reasonable, visualizing PCA-reduced embeddings from $\tilde{D}_{1}$, $\tilde{D}_{2}$, and $D_{*}$ for $\beta=0.5$ and varying $\alpha$ as training progresses For $\alpha=0.5$, the PCA shows clearly separated task clusters, indicating that both models are better tailored to the individual tasks. In contrast, the lack of well-defined clusters when $\alpha\neq 0.5$ suggests reduced sensitivity to task type, e.g. generic answers. Yet, in Table 4, we see that when $\alpha=0$, models homogenize slightly on the shared task $D_{*}$ but diverge on model-specific tasks. This makes sense since models the models do not train on each other’s tasks. Once $\alpha>0$, homogenization increases for all datasets/tasks.
@@ 7 Discussion
Limitations. Our work has several limitations. We discuss possible objections to our claims about the increasing presence of generated outputs in training datasets (e.g. (Drayson et al., 2025)) in Appendix B. Also, our theoretical framework assumes that new data in model updates is purely synthetic. In reality, if internet scrapes are used to create model update datasets, they will contain both synthetic and real data. Finally, we assume that model trainers use new scraped data for each model update but only reuse data from initial training. This assumption may limit the range of outcomes.
Broader Impacts. If data-mediated interactions homogenize generative models, causing them to coalesce on certain viewpoints, this could lead to pervasive bias in AI-generated content. Peterson (2025) discusses this possibility, while  Wenger & Kenett (2025) showed homogeneity across creative outputs from many LLMs, suggesting these homogenization effects may already be felt by models. Much future study is needed to evaluate the extent to which data-mediated interactions fuel homogeneity (as opposed to other causes) and develop mitigations.
Conclusions and Future Work. We provide a first look at possible outcomes of genAI models trained on each others’ data and find mixed effects. Training on other models’ data exposes models to concepts possibly missed in their own training data, but can homogenize model behaviors. Future work could consider additional nuances of interactions between models, explore how these interactions evolve in other modalities like image generation, and investigate whether fixed points (e.g. like the universal $\pi^{2}/6$ pathway of (Dey & Donoho, 2024)) exist under this paradigm.
@@ Appendix A Exact training data statements from LLM papers
Publicly released on HuggingFace (Anthropic, ).
GPT4+ (Achiam et al., 2023) No information provided No information provided
Grok 3 (xAI, 2025) No information provided No information provided
Jamba (Team et al., 2024b)
@@
[[Supervised fine tuning]] leverages highly
@@ Appendix B Counterarguments
Can’t we use watermarks to filter AI-generated content from future internet-scraped datasets? Several companies have publicly stated that they watermark AI-generated content (Dathathri et al., 2024; AI, 2024b; Clegg, 2024), making this argument plausible. Furthermore, (Drayson et al., 2025) show that using watermark detection techniques can help avoid model collapse under certain circumstances. However, reliance on watermarking has two major issues. First, watermarks are difficult to reliably detect and/or easily removed from generated content (Zhang et al., 2023; Longpre et al., 2024; Sadasivan et al., 2023). Second, detection via watermark requires sharing of watermark detection information, which is essentially a game-theoretic problem that relies on other companies’ willingness to cooperate. Both issues make watermarks an unreliable mechanism to rid datasets of AI-generated content.
@@
What if there’s less AI-generated online content than we think? Some work suggests there may be less AI-generated content online than previously postulated (Matatov et al., 2024). However, other works consistently point to an uptick in AI-generated content online (Sun et al., 2025). We believe the widespread adoption and use of generative AI models across industries (AI, 2024a; Handa et al., 2025; Reuters, 2024), particularly for use in content creation (gen, 2022), provides strong evidence that AI-generated content will become a regular part of online life. Further empirical work is needed to vet both claims.
@@
What if one model provider dominates online content? Although numerous generative AI models are available online, some evidence suggests that one or two companies may dominate the AI landscape. A market research firm estimated that in January 2025, Open AI’s ChatGPT had 340 million monthly active users, Microsoft Copilot had 11 million, Google Gemini had 80 million, and Anthropic’s Claude had 2 million (Zitron, 2025). If only one model/company dominates, our paradigm of training on other models’ data will no longer be relevant and collapses back to the single model setting of prior work, e.g. (Shumailov et al., 2024; Kazdan et al., 2025; Dey & Donoho, 2024). Currently, though, this market research suggests that there are several models used by millions of users, making our assumptions somewhat reasonable. Future research could analyze market trends in model use and adoption to determine realistic assumptions.
Appendix C Additional Results
Tables 6 and 7 show full results for all tested $\beta$, $\alpha$ settings for the $K=2$ setting across both the Llama and OPT model architectures. As stated in the main paper body, the $\beta=0.5$ and $\alpha=0.5$ settings perform best. Figures 45, and 6 show loss values at each generation for (1) predicted theoretical results, (2) OPT $K=2$ models, and (3) Llama $K=2$ models. Finally, Figure 7 shows results for a $K=3$ system of interacting OPT models.
Table 6: Change in loss behavior for $K=2$ interacting LLama models at various $\alpha,\beta$. We show results as $\texttt{initial}\rightarrow\texttt{final}$ prediction loss values for models on their own and the other models’ tasks, at $\texttt{initial}=T=0$ and $\texttt{final}=T=15$ generations. For clarity, we colorize loss increase, decrease, and constancy ($\Delta\leq 0.1$).
$\beta=0$ $\beta=0.5$ $\beta=1.0$
$\alpha=0$ $\alpha=0.5$ $\alpha=1.0$ $\alpha=0$ $\alpha=0.5$ $\alpha=1.0$ $\alpha=0$ $\alpha=0.5$ $\alpha=1.0$
Model 1
on Task 1
$2.8\rightarrow 4.5$ $\rightarrow 3.2 2.82.8\rightarrow$ 3.1 $2.8\rightarrow 2.9$ $2.8\rightarrow 2.9$ $\rightarrow 3.0 2.82.8$$\rightarrow 2.8$ $2.8\rightarrow 2.8$ $2.8\rightarrow 3.2$
Model 2
on Task 2
$1.2\rightarrow 2.0$ $1.2\rightarrow 1.9$ $\rightarrow 1.8 1.21.2$$\rightarrow 1.2$ $1.2\rightarrow 1.3$ $\rightarrow 1.9 1.21.2$$\rightarrow 1.2$ $1.2\rightarrow 1.2$ $1.2\rightarrow 1.8$
Model 1
on Task 2
$2.0\rightarrow 2.9$ $2.0\rightarrow 1.9$ $2.0\rightarrow 1.7$ $2.0\rightarrow 2.5$ $2.0\rightarrow 1.4$ $2.0\rightarrow 1.8$ $2.0\rightarrow 2.6$ $2.0\rightarrow 1.3$ $2.0\rightarrow 1.7$
Model 2
on Task 1
$3.7\rightarrow 5.0$ $3.7\rightarrow 3.1$ $3.7\rightarrow 3.1$ $3.7\rightarrow 4.2$ $3.7\rightarrow 3.0$ $3.7\rightarrow 3.0$ $3.7\rightarrow 4.4$ $3.7\rightarrow 2.9$ $3.7\rightarrow 3.1$
Table 7: Change in loss behavior for $K=2$ interacting OPT models at various $\alpha,\beta$ We show results as $\texttt{initial}\rightarrow\texttt{final}$ prediction loss values for models on their own and the other models’ tasks, at $\texttt{initial}=T=0$ and $\texttt{final}=T=15$ generations. For clarity, we colorize loss increase, decrease, and constancy ($\Delta\leq 0.1$).
$\beta=0$ $\beta=0.5$ $\beta=1.0$
$\alpha=0$ $\alpha=0.5$ $\alpha=1.0$ $\alpha=0$ $\alpha=0.5$ $\alpha=1.0$ $\alpha=0$ $\alpha=0.5$ $\alpha=1.0$
Model 1
on Task 1
$3\rightarrow 4.8 3.3$$\rightarrow 3.6 3.3$$\rightarrow 3.5 3.3$$\rightarrow 3.3 3.3$$\rightarrow 3.3 3.3$$\rightarrow 3.5 3.3$$\rightarrow 3.1 3.3$$\rightarrow 3.2 3.3$$3\rightarrow 3.5$
Model 2
on Task 2
$1.8\rightarrow 2.6$ $1.8\rightarrow 2.2$ $1.8\rightarrow 2.1$ $1.8\rightarrow 1.7$ $1.8\rightarrow 1.7$ $1.8\rightarrow 2.2$ $\rightarrow 1.5 1.81.8$$\rightarrow 1.8$ $1.8\rightarrow 2.2$
Model 1
on Task 2
$3.1\rightarrow 3.7$ $3.1\rightarrow 2.2$ $3.1\rightarrow 2.2$ $3.1\rightarrow 3.5$ $3.1\rightarrow 1.8$ $3.1\rightarrow 2.2$ $3.1\rightarrow 3.4$ $3.1\rightarrow 1.8$ $3.1\rightarrow 2.2$
Model 2
on Task 1
$5.1\rightarrow 5.6$ $5.1\rightarrow 3.6$ $\rightarrow 3.5 5.15.1$$\rightarrow 5.1$ $5.1\rightarrow 3.5$ $5.1\rightarrow 3.5$ $5.1\rightarrow 5.4$ $5.1\rightarrow 3.5$ $5.1\rightarrow 3.4$
@@ Appendix C Additional Results
Figure 6: Actual behavior over time for interactions between Llama models ($K=2$) with varying $\alpha,\beta$.
Figure 7: Actual behavior over time for interactions between OPT models ($K=3$) with varying $\alpha,\beta$.
Appendix D Code for experiments
@@
Appendix E Proofs for results in Section 5
@@
E.1 Proof of Theorem 1
[…]
@@ E.1 Proof of Theorem 1
For the inductive case, assume that the stated distribution holds up to generation t1t-1. From the definition of the workflow, (21) holds with $\bm{w}_{t}\sim\mathsf{N}(0,\sigma^{2}\mathrm{I})$ independent of everything else. Thus $\hat{\bm{\theta}}_{t}$ is Gaussian with mean
@@ E.2 Proof of Theorem 2
[…]
@@
Next, suppose that $\bm{G}_{1},\dots,\bm{G}_{t}$ are full rank and the stated distribution holds up to time t1t-1. By Theorem 1 we know that $\hat{\bm{\theta}}_{t}$ is Gaussian and so all that remains is to verify the given expressions for the mean and covariance. By the linearity of expectation and (21), the mean satisfies
@@ E.3 Proof of Theorem 3
E.4 Proof of Lemma 1
If $\alpha=0$ or if $\bm{S}=\bm{0}$ then $\bm{Q}=\bm{0}$ and so the stated result holds. Henceforth, we assume $0<\alpha<1$ and $\bm{S}$ is nonzero. Suppose that $\gamma\bm{S}=\lambda\tilde{\bm{S}}+(1-\lambda)(\mathrm{I}_{K}\otimes S_{*})$ for some $0<\beta\leq\lambda\leq 1$ and $\gamma>0$. Then,